40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Nov 15 17:29:14 2019, Raymond, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab IMG_0270.JPGIMG_0269.JPG
    Reply  Tue Nov 19 20:16:40 2019, Raymond, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab IMG_0483.JPG
       Reply  Wed Nov 20 11:59:17 2019, Koji, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab 
          Reply  Wed Nov 20 12:25:58 2019, Koji, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab 
             Reply  Wed Nov 20 17:44:10 2019, Koji, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab P_20191120_174034_vHDR_On.jpgP_20191120_174040_vHDR_On.jpg
                Reply  Thu Nov 21 21:45:10 2019, Raymond, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab DT02H_18_32PN.pdf
                   Reply  Fri Nov 22 13:21:26 2019, Raymond, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab IMG_0497.JPGIMG_0498.JPG
                      Reply  Fri Nov 22 15:33:40 2019, Koji, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab P_20191122_150845_vHDR_On.jpg
                         Reply  Sat Nov 23 13:00:06 2019, Koji, Lab Infrastructure, Optics, Optical window transmission measurement optical_window_T.pdfP_20191122_183426_vHDR_On.jpgP_20191122_183436_vHDR_On.jpg
                         Reply  Tue Nov 26 15:36:59 2019, Raymond, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab IRLabsQIL.pdf
                            Reply  Wed Dec 4 13:54:04 2019, Raymond, Lab Infrastructure, General, Moved IR Labs cryostat into QIL from Cryo Lab 
Message ID: 2458     Entry time: Sat Nov 23 13:00:06 2019     In reply to: 2457
Author: Koji 
Type: Lab Infrastructure 
Category: Optics 
Subject: Optical window transmission measurement 

The power transmission of the optical window for the IRLab cryostat was measured to be 0.966+/-0.002 at 2004nm. (Attachment 1)

A chopper powermeter was set to the QE measurement setup (Attachment 2). The window was held with a mount as shown in Attachmnent 3. The laser source was excited with the pumping current of 101.04mA. The output power was  monitored with a Thorlabs DET10D (PD#2 with Amp#2) attached at the 10% side of the 90:10 beamsplitter. The detected photocurrent after subtracting the dark current of 15.7uA was 152uA. The power meter detected the power around 0.95mW, while the power with the window inserted was around 0.91~0.92.


PD1       Window   No Window
[V]       [mW]     [mW]
-0.855    0.913    0.944
-0.855    0.906    0.951
-0.855    0.914    0.947
-0.855    0.922    0.950
-0.855    0.913    0.949
-0.855    0.912    0.948
-0.855    0.920    0.946
-0.855    0.915    0.946
-0.855    0.916    0.951
-0.855    0.915    0.952
-0.855    0.919    0.947
-0.855    0.921    0.944
-0.855    0.916    0.948

Note: PD1 had the dark output of -0.0809V.
Note2: The power meter readings had the fluctuation of +/-0.005 mW

Attachment 1: optical_window_T.pdf  16 kB  | Hide | Hide all
optical_window_T.pdf
Attachment 2: P_20191122_183426_vHDR_On.jpg  888 kB  | Hide | Hide all
P_20191122_183426_vHDR_On.jpg
Attachment 3: P_20191122_183436_vHDR_On.jpg  1.245 MB  | Hide | Hide all
P_20191122_183436_vHDR_On.jpg
ELOG V3.1.3-